OpenStudy (awesomedude2078):

Loan Amount: $13,000 Method of payment: discount basis Amount of Interest: $625 Term of loan: 1 year Effective Rate of Interest (to the nearest tenth) = _____.

OpenStudy (awesomedude2078):

@tkhunny

OpenStudy (tkhunny):

These are often done incorrectly. I'll show you one for free. Simple Discount is to extract the Loan Interest from the Beginning of the loan period. 625 / 13000 = 4.8076923% Simple Discount Rate 13000 - 625 = 12375 = Cash in Your Pocket at Loan Inception 625 / 12375 = 5.0505051% Equivalent Simple Interest Rate Your turn. You do one.

OpenStudy (awesomedude2078):

thanks

OpenStudy (tkhunny):

I'm looking for your work...